Inkeq Page 1 > of 3 1/T (1/K) 3) For the binding of Protein X to Protein Y, where kon = 1EMs and kom = 1E s. a. What is th
Answer #1

KD is the measure of protein binding affinity and is given by Kp koff kon

a.Affinity of protein X for protein Y, 15-- Kp = 15 M-1,-1 = 1E-7M

b. For [Y]=1nM=1 E-9M, associate rate= kon [X] [Y] = kon Y2 = 1EM-1:-1 (1E-9M)2 = 1E-13M 5-1

dissociate rate= koff [XY] = koff [Y] = 1E--+ x 16-18 M = 15-20 M45-

for [Y]=100nM=1E-7M,

associate rate= kon [X] [Y] = kon [Y]° = 1EⓇM-1*(1E-7M) = 1E-9M5-1

dissociate rate= kof{ [XY] = koji [72] = 1E--1 x (1E-?M? = 1E-16 Mºs-1

for [Y]=10\muM=1E-6M

associate rate= kon Y 2 = 1EM-1:-1 (1E-6M)2 = 1E-?Ms-1

dissociate rate= koff [Y2 = 15-25-1 x 1E-12M = 1E-14 Ms-1